### v2

# All-electrically controlled spintronics in altermagnetic heterostructures

Abstract
The recent development of altermagnetic materials, supporting spin splitting without net magnetization, opens new directions for spintronics that are fundamentally distinct from conventional ferromagnetic, antiferromagnetic, or spin-orbit coupling systems. Here we investigate spin-selective quantum transport in heterostructures composed of a normal metal and a two-dimensional $d$-wave altermagnet. We focus on two types of $d$-wave altermagnets, namely, weak and strong altermagnets that support close elliptic and open hyperbolic spin-resolved Fermi surfaces, respectively. Building on these distinct electronic structures, we propose all-electrically controlled spin filter and spin valve devices, where quantum resonant tunneling enables highly spin-polarized conductance tunable via gate voltage and interface transparency. In particular, we find that strong altermagnets allow gate-tunable full spin polarization that is robust against interface scattering and can be reversed by gate control. We further demonstrate that a double-gated spin valve electrically switches between parallel and antiparallel spin configurations, analogous to magnetic junctions but without the need for external magnetic fields. Our results establish both weak and strong altermagnets as promising platforms for realizing magnetic-field-free electrically tunable spintronic functionalities.
I Introduction
Unconventional magnets have emerged as a new class of systems that lie beyond the traditional dichotomy of ferromagnets and antiferromagnets [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. These so-called third type of magnets support spin-split Fermi surfaces resembling those of ferromagnets [12, 13], yet exhibiting globally vanishing magnetization due to a compensated magnetic ordering, akin to antiferromagnets [14, 15, 16]. Depending on the momentum-space symmetry of the exchange field and the underlying crystalline symmetries, unconventional magnets can be classified according to the angular character of the spin splitting, including $p$-, $d$-, $f$-, $g$-, and $i$-wave types [7, 8, 9, 17, 18]. Among these, $d$-, $g$-, and $i$-wave magnets are known as altermagnets (AMs), characterized by simultaneously broken time-reversal symmetry and spatial rotational symmetry, but preserving their joint symmetry [7, 8]. Furthermore, weak and strong AMs can be distinguished by the geometry of their closed and open Fermi surfaces, respectively. [7, 8, 19, 20, 21]. In contrast, $p$- and $f$-wave magnets break only rotational symmetry but preserve time-reversal symmetry originating from nonsymmorphic lattice symmetries [9, 10], typically exhibiting closed anisotropic Fermi surfaces.
Unconventional magnets are promising platforms for spintronic devices [13, 14]. The anisotropic spin split Fermi surface offers mechanisms for giant magnetoresistance [22], spin-orbit-free anomalous Hall effects [23, 1, 24, 25, 26], spin-transfer torques [27, 28, 29], spin filtering effects [30], spin pumping effects [31], light-matter interactions [32, 33], light-induced spin density [34], non-Hermitian electronic responses [35, 36], strongly correlation in Mott insulators [37] and other novel phenomena [38, 39]. These mechanisms are expected in experimentally accessible altermagnetic materials such as RuO2 [40, 23, 25], MnTe [41, 24], Mn5Si3 [26, 42], CrSb [43], and Mn2Au [44], exhibiting novel spin transport. In addition to their normal-state properties, unconventional magnets have recently garnered attention in the context of superconducting spintronics by inducting high-parities spin-triplet states [34, 45, 46, 47, 48, 49], orientation-dependent transport in altermagnetic-supercondctor hybridstructure [50, 51, 21, 52, 53, 54, 20, 21, 55], gate-controlled Josephson junctions [46, 56, 57, 58, 59, 60], topological superconductors [61, 49, 62, 63, 64, 65], nonreciprocal field-free superconducting devices [66, 67, 68] and ther superconducting phenomena generated by the interplay between unconventional magnetism and superconductivity [11]. Together, these developments establish unconventional magnets as a fertile ground for exploring spintronic functionalities without net magnetization.
At the moment, there exist several proposals for spintronic devices based on AMs [5, 6, 11]. For instance, spin-dependent tunneling barriers have been shown to induce spin splitting in insulating AMs [30]. Alternatively, $g$-wave altermagnetic semiconductors have demonstrated strain-induced spin-orbit coupling, enabling gate-controlled spin splitting [69]. More recently, electrically gated spin-layer coupling can be achieved by applying different potentials to individual layers [70]. In recently developed ferroelectric-switchable AMs [71, 72], the spin-split Fermi surface can be reoriented by reversing the ferroelectric polarization. Despite these proposals offering a pathway toward AM-based spintronics, the unique non-relativistic nature of spin splitting in AM systems remains underutilized and largely unexplored. Furthermore, most theoretical and experimental efforts have focused on weak AMs, while the transport properties of strong AMs, featuring open Fermi surfaces, remain largely underexplored [19, 20, 21]. Thus, leveraging the intrinsic non-relativistic spin splitting of AMs for realizing spintronic functionalities without external magnetic fields remains an open and compelling challenge.
Figure 1: Schematic of electrically controlled spintronic devices based on AM heterostructures: (a) spin filter and (b) spin valve. (a) In the spin filter, a gate applied to the AM region generates fully spin-polarized conductance by selectively blocking one spin channel. (b) The spin valve consists of two gated spin filters, exhibiting electrically tunable switching between high and low conductance states. This behavior is analogous to the parallel and antiparallel magnetization configurations in conventional ferromagnetic bilayer spin valves, but achieved here without magnetic fields and net magnetization.
In this work, we explore spin-selective transport in heterostructures composed of AMs and normal metals, aiming to establish all-electrical spintronic functionality that harnesses the intrinsic, nonrelativistic spin splitting of AMs. Two AM-based spintronics devices with fundamental interest are proposed, including a spin filter and spin valve (Fig. 1). In the regime of quantum-coherent transport, a conventional spin filter is realized in a ferromagnet junction, where the Zeeman splitting in combination with a build-up potential barrier gives rise to a spin-polarized flow since spin-up and spin-down electrons experience different barrier heights [73, 12, 13]. While the spin polarization of the current through the junction is determined by the magnetization of the ferromagnet, the applied barrier can tune the allowed occupation of states belonging to the lower Zeeman level and thus strengthen the spin filtering. In the proposed spin filter [Fig. 1(a)], a gate applied to the AM region generates switchable fully spin-polarized conductance by selectively blocking one spin channel, yet without requiring a Zeeman exchange field. Furthermore, two AM-based spin filters connected in series compose a spin valve [Fig. 1(b)]. This is similar to the conventional spin valve composed of a conventional magnetic bilayer, where only electrons with spin aligned to the magnetization contribute to the current [13, 12]. Consequently, conductance is high when the magnetization directions of the two ferromagnetic layers are aligned, while it is suppressed when these directions are opposite, known as the parallel and antiparallel configurations, respectively. Parallel-antiparallel switching is typically achieved by applying an external magnetic field, which controls the relative magnetization orientations between layers. The resulting difference in conductance and thus the resistance under field reversal gives rise to the giant magnetoresistance effect [74, 75], which forms the basis for modern electronics applications such as data storage technologies, magnetic random access memory, and magnetic field sensors [12, 13, 14, 15, 16]. The behaviors analogous to the parallel-antiparallel switching are also found in the proposed spin valve based on strong AMs, where the alteration of the spin polarization in each spin filter is achieved by the applied gate [Fig. 1(b)], without an external magnetic field and net magnetization.
To elaborate on the mechanism of these devices, we begin by classifying weak and strong AMs based on the competition between the spin-dependent exchange interaction and kinetic energy (Fig. 2). We find that their distinct Fermi surfaces give rise to qualitatively different transport characteristics when assembled into junctions (Fig. 3). In particular, quantum resonant tunneling leads to spin-selective conductance that can be controlled by a gate voltage and interface transparency. For weak AMs, spin filtering arises through barrier-enhanced conductance polarization due to the suppression of one spin species by the confinement effect [76, 77, 78, 79]. In contrast, strong AMs enable fully spin-polarized conductance over wide energy ranges due to the intrinsic absence of one spin species near the Fermi level. We find that this intrinsic spin-filtering effect is robust against interface scattering and does not require magnetic fields or spin-orbit coupling (Fig. 4). Based on this mechanism, we further propose a double-gated spin valve device in a strong AM (Fig. 5). Interestingly, in this device, alternating between high and low conductance states can be realized by electrical gating, analogous to parallel and antiparallel spin configurations in conventional magnetic spin valves [13, 14] but without requiring net magnetization or an external magnetic field. Our results reveal that the non-relativistic spin splitting of AM offers a promising ground for engineering robust and tunable spintronic devices based entirely on electric field control. The proposed all-gate-controlled altermagnetic spin filter and spin valve are enabled by the intrinsic Fermi surface structure of weak and strong altermagnetic materials, without the need for net magnetization, relativistic spin-orbit coupling, or external magnetic fields.
The remainder of this work is organized as follows. The band structure, Fermi surface features of weak and strong AM are demonstrated in Sec. II, which determines the transport properties and the spin filter effect in the altermagnetic junction illustrated in Sec. III. Based on these properties, two spintronic devices, gate-controlled spin filter and spin valve, are proposed in Sec. IV. The conclusion is given in Sec. V.
II Weak and strong altermagnets
We start by introducing the Hamiltonian of $d$-wave AMs in the spin basis $\left(\psi_{\bm{k},\uparrow},\psi_{\bm{k},\downarrow}\right)^{\mathrm{T}}$ [7, 8]
$H_{\text{AM}}(\bm{k})=\left(\begin{array}[]{cc}H_{+}(\bm{k})&0\\ 0&H_{-}(\bm{k})\end{array}\right)\text{,}$ (1)
where
$H_{\sigma}(\bm{k})=\xi_{\bm{k}}+J_{\bm{k}},$ (2)
is the spin-resolved Hamiltonian for electrons with spin $\sigma=+1$ ($\sigma=-1$) being parallel (antiparallel) to the Néel vector, which is hereafter considered along $z$. The kinetic energy $\xi_{\bm{k}}$ and altermagnetic exchange field $J_{\bm{k}}$ are defined, respectively, as
$\xi_{\bm{k}}$ $=$ $ta^{2}\left(k_{x}^{2}+k_{y}^{2}\right)\text{,}$ (3)
$J_{\bm{k}}$ $=$ $\sigma Ja^{2}\left[\left(k_{x}^{2}-k_{y}^{2}\right)\cos 2\theta_{J}+2k_{x}k_{y}\sin 2\theta_{J}\right]\text{.}$ (4)
Here, $t$ is the nearest-neighbor hopping energy (set as the energy unit $t=1$), and $a$ is the lattice constant (set as $a=1$). The two-dimensional momentum is denoted by $\bm{k}=(k_{x},k_{y})$, with magnitude $k=|\bm{k}|$ and polar angle $\phi_{k}=\arctan(k_{y}/k_{x})$. The parameter $J$ represents the strength of the altermagnetic field, and $\theta_{J}$ defines the angle between the altermagnetic lobe and the $x$-axis.
The momentum-dependent spin splitting is captured by Eq. (4), which describes a $d_{x^{2}-y^{2}}$-wave ($d_{xy}$-wave) AM when $\theta_{J}=0$ ($\theta_{J}=\pi/4$[7, 8], hosting an anisotropic spin split $d$-wave Fermi surface that depends on the momentum direction $\phi_{k}$ for a given $\theta_{J}$. The maximum altermagnetic effect occurs along momentum directions $\phi_{k}=\theta_{J}+n\pi/2$, with $n\in\mathbb{Z}$, whereas it vanishes along directions $\phi_{k}=\theta_{J}+(2n+1)\pi/4$.
The competition between the altermagnetic term, Eq. (4), and the kinetic energy term, Eq. (3), defines weak and strong altermagnetism. To illustrate this effect, we first consider the $d_{x^{2}-y^{2}}$-wave AM as an example, whose dispersion can be obtained from Eq.(2) with $\theta_{J}=0$, as
$E_{\sigma}=\left(t+\sigma J\right)a^{2}k_{x}^{2}+\left(t-\sigma J\right)a^{2}k_{y}^{2}\text{.}$ (5)
Figure 2 displays $E_{\sigma}$ [Eq. (5)] as a function of $k_{x,y}$. There are two directions of maximal altermagnetic effect, namely along the $k_{x}$ and $k_{y}$ axes for $k_{y}=0$ and $k_{x}=0$, respectively. Along these directions, the dispersion of a weak AM ($J<t$) [20] consists of two upward-opening parabolas with spin- and direction-dependent curvature [Fig. 2(a)]. For a given energy $E$, the Fermi surface satisfies the equation of a standard ellipse,
$\frac{a^{2}k_{x}^{2}}{E/\left(t+\sigma J\right)}+\frac{a^{2}k_{y}^{2}}{E/\left(t-\sigma J\right)}=1\text{,}$ (6)
which describes two spin-dependent elliptical Fermi surfaces with mutually orthogonal principal axes, with semi-major and semi-minor axes being $\sqrt{E/(t+\sigma J)}/a$ and $\sqrt{E/(t-\sigma J)}/a$ depending on spin [see the inset of Fig. 2(a)].
Figure 2: Dispersion of (a) weak and (b) strong AM along $k_{x}$ and $k_{y}$ directions with $k_{y}=0$ and $k_{x}=0$, respectively. The insets exhibited the anisotropic Fermi surface at $E/t=1$ for weak altermagnet and $E/t=\pm 1$ for strong altermagnet as indicated by the horizontal dashed line. The spin-up (spin-down) subband is in blue (red). The altermagnetic strength is (a) $J/t=0.5$ and (b) $J/t=1.5$. Parameters: $t=1$ is the energy unit and $a=1$ is the lattice constant.
In contrast, for a strong AM ($J>t$) [20], the dispersion consists of two parabolic bands with opposite opening directions and spin-dependent curvature [see Fig. 2(b)]. This arises because the sign of the effective mass, determined by the coefficients $(t\pm\sigma J)a^{2}$ in Eq. (5), changes along orthogonal momentum directions. As a result, the band structure hosts a saddle point [80] at the spin-degenerate momentum $(k_{x},k_{y})=(0,0)$, where the curvature is positive in one direction and negative in the perpendicular one. Such a saddle point gives rise to a logarithmic van Hove singularity in the density of states near the corresponding energy [80]. The opposite opening directions of the spin-dependent bands allow electronic states to exist below $E=0$, with spin-up and spin-down states shifted oppositely in energy [Fig. 2(b)]. This energy asymmetry arises from the spin-dependent curvature of the bands in orthogonal directions. As a result, the strong AM is characterized by open, spin-dependent anisotropic Fermi surfaces, which take the form of two orthogonally oriented sets of hyperbolas. These Fermi contours, distinct from the closed elliptical shapes found in the weak AM case, are described by
$\sigma\frac{a^{2}k_{x}^{2}}{E/\left(\sigma t+J\right)}-\sigma\frac{a^{2}k_{y}^{2}}{E/\left(J-\sigma t\right)}=1\text{,}$ (7)
for a fixed energy $E$. These open Fermi surfaces impose directional constraints on the availability of spin-polarized electronic states. As shown in the inset of Fig. 2(b), for a given positive energy $E$, spin-up (spin-down) states are absent when $a|k_{x}|<\sqrt{E/(t+J)}$ [$a|k_{y}|<\sqrt{E/(t+J)}$], due to the absence of real solutions for the corresponding hyperbolic Fermi contour. Conversely, for negative energies $E<0$, spin-up (spin-down) states are absent when $a|k_{y}|<\sqrt{|E|/(t-J)}$ [$a|k_{x}|<\sqrt{|E|/(t-J)}$]. This directional open Fermi surface is a hallmark of the strong altermagnetic regime and stands in sharp contrast to the weak AM case, where both spin species coexist over closed elliptical Fermi surfaces. While our discussions above are based on $d_{x^{2}-y^{2}}$-wave AM with $\theta_{J}=0$, the distinct properties such as the dispersions, Fermi surface geometry between weak and strong AM, are shared in all $d$-wave AM with arbitrary $\theta_{J}$.
III Spin-Selective Transport in Altermagnetic Junctions
The contrasting spin-split Fermi surfaces in weak and strong AMs give rise to qualitatively distinct transport responses for designing spintronic devices. A central objective in spintronics is generating and manipulating spin-polarized conductance or current, particularly through electrically tunable spin filters [12, 13, 14, 15, 16]. To explore this device, we consider a junction composed of an AM attached with two normal metal leads, as schematically shown in Fig. 3(a-i). This heterostructure supports a gate-controllable spin-filtering effect, enabled by the intrinsic energy- and momentum-dependent spin splitting in the altermagnetism. Below, we show spin-polarized transport by using a scattering approach.
III.1 Scattering formalism and spin-polarized conductance
We consider a spin filter junction oriented along $x$ direction, as shown in Fig. 3(a-i). This device is composed of a gated AM sandwiched between two normal metal leads. The spin filter can be controlled by the gate $U_{g}$ applied in the AM region and the tunneling barrier at the interfaces between the normal leads and the AM. The Hamiltonian below models the junction
$H_{\text{SF}}^{\sigma}\left(x\right)$ $=$ $ta^{2}k_{y}^{2}+ta^{2}\left(-i\partial_{x}\right)^{2}-U(x)$
$-\sigma J(x)a^{2}k_{y}^{2}\cos\left(2\theta_{J}\right)$
$+a^{2}k_{y}\sin\left(2\theta_{J}\right)\sigma\left\{J(x),-i\partial_{x}\right\}$
$+a^{2}\cos\left(2\theta_{J}\right)\sigma\partial_{x}J(x)\partial_{x},$
where
$U(x)$ $=$ $U_{L}\left[\Theta(-x)+\Theta(x-d)\right]-U_{g}\Theta(x)\Theta(d-x)$ (9)
$-V(\delta_{x,0}+\delta_{x,d})\text{,}$
describes the electrostatic gating potential applied across the junction, with $\Theta(x)$ the Heaviside step function and $\delta_{x,x_{i}}$ the Kronecker delta. Here, $U_{L}$ sets the chemical potential in the left ($x<0$) and right ($x>d$) normal leads, $U_{g}$ denotes the gate voltage applied to the central AM region, and $V$ represents the tunneling barriers at the normal–AM interfaces located at $x=0$ and $x=d$. Here, we assume the chemical potentials in both leads are identical, and the same tunneling barriers are applied at the two interfaces. The AM occupies the central region $0<x<d$, with
$J(x)=J\,\Theta(x)\Theta(d-x),$ (10)
[…]
The transport properties of the junction can be obtained by solving the quantum scattering problem [76, 81]. For a spin-$\sigma$ incident electron with energy $E$ in the direction $\theta_{k}$, the eigenfunction is given by $\Psi^{\sigma}(x)=\psi^{\sigma}(x)e^{ik_{y}y}$, with
$\psi^{\sigma}(x)=\left\{\begin{array}[]{ll}e^{ik_{0}x}+r_{\sigma}e^{-ik_{0}x},&x\leq 0,\\[4.0pt] c_{1}e^{ik_{+}^{\sigma}x}+c_{2}e^{ik_{-}^{\sigma}x},&0\leq x\leq d,\\[4.0pt] \tau_{\sigma}e^{ik_{0}x},&x\geq d,\end{array}\right.$ (11)
where $r_{\sigma}$ and $\tau_{\sigma}$ are the reflection and transmission coefficients, respectively, and $c_{1,2}$ are the scattering amplitudes in the central AM region. The longitudinal and transverse momentum are defined as
$k_{0}=k_{F}\cos\theta_{k}\text{,}$ (12)
and
$k_{y}=k_{F}\sin\theta_{k}\text{,}$ (13)
in the normal lead region, with $k_{F}=\sqrt{(E+U_{L})/(ta^{2})}$ being the Fermi wave number of the incident electron. The electron injection angle satisfies
$\theta_{k}=\arctan(k_{y}/k_{0})\text{.}$ (14)
The longitudinal in the AM region is
$k_{\pm}^{\sigma}=\frac{-\sigma Jak_{y}\sin(2\theta_{J})\pm\Omega}{at_{\sigma}}\text{,}$ (15)
with $\Omega=\sqrt{(J^{2}-t^{2})a^{2}k_{y}^{2}+t_{\sigma}(E-U_{g})}\text{,}$ and $t_{\sigma}=t+\sigma J\cos(2\theta_{J})$, which depends on the strength $J$, orientation $\theta_{J}$, and the gate $U_{g}$ applied in AM, as well as the incident angle $\theta_{k}$ and energy $E$ of the injecting electron. Using the boundary conditions associated with Eq. (III.1), we impose
$\Psi^{\sigma}(0^{-})$ $=$ $\Psi^{\sigma}(0^{+}),$
$\Psi^{\sigma}(d^{-})$ $=$ $\Psi^{\sigma}(d^{+}),$ (16)
$-t_{-\sigma}\partial_{x}\Psi^{\sigma}(0^{+})+t\partial_{x}\Psi^{\sigma}(0^{-})$ $=$ $-V_{\sigma}\Psi^{\sigma}(0),$
$t_{-\sigma}\partial_{x}\Psi^{\sigma}(d^{-})-t\partial_{x}\Psi^{\sigma}(d^{+})$ $=$ $-V_{\sigma}^{*}\Psi^{\sigma}(d),$
where $t_{-\sigma}=t-\sigma J\cos(2\theta_{J})$ and $V_{\sigma}=V/a^{2}-\sigma iJk_{y}\sin(2\theta_{J})$. By substituting the eigenfunction [Eq. (11)] into the boundary condition [Eq. (16)], the spin-resolved transmission amplitude is then given by
$\tau_{\sigma}(E,\theta_{k})=2k_{0}\frac{(t_{\sigma}/t)(k_{+}^{\sigma}-k_{-}^{\sigma})e^{-id(k_{0}-k_{+}^{\sigma}-k_{-}^{\sigma})}}{e^{idk_{-}^{\sigma}}Z^{+}_{+}Z^{-}_{-}-e^{idk_{+}^{\sigma}}Z^{-}_{+}Z^{+}_{-}},$ (17)
with
$Z^{\alpha}_{\pm}=k_{0}+iV/(ta^{2})+\alpha\frac{t_{-\sigma}k_{\pm}^{\sigma}+\sigma Jk_{y}\sin 2\theta_{J}}{t},$ (18)
and $\alpha=\pm 1$. The resulting spin-resolved transmission probability is
$T_{\sigma}(E,\theta_{k})=\left|\tau_{\sigma}(E,\theta_{k})\right|^{2},$ (19)
which determines the spin-resolved conductance at zero temperature [82]
$G_{\sigma}(E_{F})=G_{0}\int_{-\pi/2}^{\pi/2}T_{\sigma}(E,\theta_{k})\cos\theta_{k}\,d\theta_{k},$ (20)
where $G_{0}=e^{2}Wk_{F}/(2\pi h)$ is the conductance unit, and $W$ is the sample width. The total conductance of the junction is given by the sum over spin channels,
$G=G_{\uparrow}+G_{\downarrow},$ (21)
where the labels $\uparrow$ and $\downarrow$ correspond to $\sigma=+1$ and $\sigma=-1$, respectively. In the linear region, the spin-resolved current driven by an applied bias $V_{\text{bias}}$ is $I_{\sigma}=G_{\sigma}V_{\text{bias}}$ and total current is $I=I_{\uparrow}+I_{\downarrow}$ [82]. To quantify the contribution of each spin channel, we define the spin polarization as
$P=\frac{G_{\uparrow}-G_{\downarrow}}{G_{\uparrow}+G_{\downarrow}},$ (22)
where $P>0$ ($P<0$) indicates a dominant contribution from spin-up (spin-down) electrons. The spin polarization $P$ serves as a measure of spin-filtering efficiency when a spin-degenerate electron beam is injected. In particular, $P=+1$ ($P=-1$) corresponds to a fully spin-polarized conductance and current carried entirely by spin-up (spin-down) states. The generation and electrical control of spin-polarized conductance and thus the current are essential goals in the development of spintronic devices.
Figure 3: (a) Schematic illustration of the spin-filter effect. (a-i) The spin-filter is composed of normal metals and a gated altermagnet by $U_{g}$. The junction is along the $x$-direction, keeping periodic boundary conditions along the $y$-direction. (a-ii): The normal-metal leads exhibit spin-degenerate parabolic dispersions (black lines), while the weak (strong) AM features spin-dependent parabolic bands with the same (opposite) curvature direction, with blue and red lines denoting the spin-up and spin-down states, respectively. The dispersions are exhibited along $k_{x}$ direction with $k_{y}=0$. (a-iii) and (a-iv): Isotropic Fermi surface of normal-metal leads and anisotropic closed and open spin-resolved Fermi surfaces for weak and strong AMs, respectively. The Fermi surfaces are exhibited in $k_{x}$-$k_{y}$ plane. Blue (red) shaded regions indicate transverse mode ($k_{y}$) ranges forbidden for spin-up (spin-down) electrons due to the AM band structure. (b) Transmission probability, spin-resolved conductance and spin polarization for a weak AM. (b-i) and (b-iii): Energy and incident-angle ($\theta_{k}$) resolved transmission probabilities $T_{\uparrow}$ and $T_{\downarrow}$. (b-ii) and (b-iv): Transmission for normal incidence ($\theta_{k}=0$) at various tunneling barrier heights $V$. The horizontal dashed line denotes the resonant energy levels [Eq. (24)] due to the confinement condition [Eq. 23]. (b-v) and (b-vi): Energy-dependent spin-resolved conductance and resulting spin polarization. (c) Same as panel (b), but for a strong AM. Parameters: $U_{L}=0$, $U_{g}=1$, $V=0$ in (i, iii, v), and junction length $d=20a$. The $d_{x^{2}-y^{2}}$-wave AM with $\theta_{J}=0$ is considered.
III.2 Transport in weak altermagnetic junctions
Based on the scattering formalism above, in this section, we analyze transport through a spin filter based on a weak $d_{x^{2}-y^{2}}$ [Fig. 3(a-ii) and (a-iii)], characterized by a closed, spin-dependent Fermi surface [Fig. 2(a)]. Figure 3(b) presents the energy- and angle-resolved transmission probabilities, as defined in Eq. (19), along with the resulting spin-resolved conductance [Eq. (20)] and spin polarization [Eq. (22)].
The spin-dependent transmission probabilities $T_{\uparrow}$ and $T_{\downarrow}$ are shown in Figs. 3(b-i) and 3(b-iii), respectively, as functions of the incident energy $E$ and injection angle $\theta_{k}$. The corresponding normal-incidence transmission profiles, $T_{\uparrow}(\theta_{k}=0)$ and $T_{\downarrow}(\theta_{k}=0)$, are plotted in Figs. 3(b-ii) and 3(b-iv) for different values of the tunneling barrier height $V$. Resonant transmission ($T_{\sigma}=1$) is achieved when the junction length $d$ satisfies a standing wave condition [76] such that it equals an integer multiple of half the wavelength of the spin-$\sigma$ electron in the altermagnetic region, where $\lambda_{\sigma}=2\pi/k_{+}^{\sigma}$. This results in the resonance condition,
$k_{+}^{\sigma}d=n\pi,\quad(n=1,2,3,\cdots),$ (23)
which ensures constructive interference between the left-going and right-going propagating modes in AM in Fabry–Pérot–like structures [76, 77, 78, 79]. So that the phase difference accumulated over distance $d$ between the two modes is an integer multiple of $\pi$. With this condition, one can check that the denominator of Eq. (17) becomes minimal, resulting in maximal transmission probability. By substituting Eq. (15) into the resonance condition [Eq. (23)], the corresponding resonant energy levels related to the confined standing wave are
$E_{n}^{\sigma}(\theta_{k})=t\frac{t_{+}(n\pi a/d)^{2}+U_{g}+t_{-}\sin^{2}\theta_{k}\,U_{L}/t}{t-t_{-}\sin^{2}\theta_{k}},$ (24)
with $t_{\pm}=t\pm\sigma J$. The resonant energy levels depend on the spin index $\sigma$, the transverse incident angle $\theta_{k}$, and device parameters such as $U_{L}$, $U_{g}$, and $d$, but are independent of the tunneling barrier $V$. Thus, the transmission probability remains unity even in the presence of a finite tunneling barrier $V$ when the incident energy satisfies the resonance condition, $E=E_{n}^{\sigma}(\theta_{k})$, but becomes strongly suppressed in the off-resonant regime, $E\in(E_{n}^{\sigma},E_{n+1}^{\sigma})$ [See Fig. 3(b-i)-(b-iii) for $n=1$]. The energy spacing between neighboring resonant levels is
$\delta_{n,E}^{\sigma}(\theta_{k})$ $=$ $E_{n+1}^{\sigma}(\theta_{k})-E_{n}^{\sigma}(\theta_{k})$
$=$ $ta^{2}\pi^{2}\frac{t+\sigma J}{t-(t-\sigma J)\sin^{2}\theta_{k}}\frac{2n+1}{d^{2}},$
which is spin-dependent. The level spacing decreases with increasing junction length $d$, but increases with the incident angle $\theta_{k}$. Since $\delta_{E}^{+}(\theta_{k})>\delta_{E}^{-}(\theta_{k})$, finite $T_{\downarrow}$ may appear in energy regions where $T_{\uparrow}$ is suppressed [See Fig. 3(b-i)-(b-iii)]. This asymmetry spin channel leads to spin-polarized tunneling. The degree of suppression increases with $V$, as the barrier height reduces the amplitude of evanescent modes within the altermagnetic region and enhances the energy selectivity of the resonance, thereby narrowing the transmission windows and increasing the contrast between on- and off-resonant transport. This behavior is illustrated in Fig. 3(b-ii) and Fig. 3(b-iv), which show the normal-incidence transmission probabilities $T_{\uparrow}(\theta_{k}=0)$ and $T_{\downarrow}(\theta_{k}=0)$, respectively, for various values of the tunneling barrier $V$. As $V$ increases, transmission away from resonance is progressively suppressed, while the resonant peaks remain robust, highlighting the role of $V$ in enhancing spin-dependent energy filtering. In the normal incident case, the lowest resonant peak occurs at $E=E_{1}^{\sigma}(0)=(t+\sigma J)(\pi a/d)^{2}+U_{g}\approx U_{g}$ for long junctions ($d\gg a$) and tunneling for $E<U_{g}$ is negligible since the incident electron energy lies below the bottom of the spin-dependent parabolic dispersion set by $U_{g}$ [Fig. 2(a-ii)]. For general injection angles, Eq. (24) implies that only electrons within a restricted angular range contribute to the transmission. This angular window is bounded by a critical angle $\theta_{c}^{\sigma}$ for spin $\sigma$, given by
$\theta_{c}^{\sigma}=\arcsin\sqrt{\frac{E-U_{g}}{\left(1-\sigma J/t\right)(E+U_{L})}}.$ (26)
Only electrons with incident angle $\theta_{k}^{\sigma}<\theta_{c}^{\sigma}$ can contribute to transmission. This spin-dependent critical angle arises from the anisotropic spin-split Fermi surface of AMs. For instance, in a $d_{x^{2}-y^{2}}$ AM with $\theta_{J}=0$, and a junction oriented along the $x$-direction [Fig. 3(a-iii)], spin-up electrons exhibit a broader angular transmission window compared to spin-down electrons.
By summing over all possible incident directions, the spin-resolved conductance [Fig. 3(b-v)] exhibits the spin-dependent resonance condition described by Eq. (23). Specifically, the spin-up conductance is significantly suppressed by the tunneling barrier except for the resonant peaks, while the spin-down conductance remains measurable. These resonant features correspond to constructive interference conditions for the spin-dependent wavefunctions inside the altermagnetic region. As the tunneling barrier height $V$ increases, the off-resonant transmission is further suppressed, thereby enhancing the spin polarization $P$ [Fig. 3(b-vi)]. However, because of the continuous and overlapping band structure of the weak AM, both spin channels inevitably contribute to transport for a given energy, see Fig. 3(b-i) and (b-iii). Electrons incident at oblique angles ($\theta_{k}\neq 0$) retain finite tunneling probabilities. This leads to a residual spin-up contribution that limits the maximum achievable spin polarization. Consequently, even in the presence of a high tunneling barrier, full spin polarization ($P=1$ or $-1$) is hardly achieved. Instead, the spin polarization saturates at approximately $P\sim 80\%$ for large $V$, as demonstrated in Fig. 3(b-vi).
The partial polarization in weak AM discussed above poses two notable challenges for device applications: (i) The maximum spin polarization is not fully tunable and remains sensitive to the Fermi energy $E$. This limits the flexibility of weak AM-based spin filters compared to conventional ferromagnetic systems [12, 13], where full spin polarization can be realized and externally modulated through magnetic fields [83] or spin-transfer torques [84]. (ii) Enhancing spin filtering requires a high tunneling barrier, which inherently reduces the total conductance and the current. The inverse relationship between spin polarization and conductance imposes a fundamental constraint on device performance, especially in regimes where a sizable spin-polarized current is required for detection or functionality. While the challenge (i) can be mitigated by applying gate voltages to shift the resonance conditions as demonstrated in Sec. IV.1, the challenge (ii) is intrinsic to the weak AM band structure. However, as we show in the following, this limitation can be circumvented by employing strong AM materials. In that case, the emergence of spin-polarized transport arises from the intrinsic band structure of the strong AM itself, rather than relying on interfacial tunneling effects. This results in robust and fully spin-polarized transport, even in the absence of a tunneling barrier.
III.3 Transport properties in strong altermagnetic Junctions
The distinct dispersion of a strong AM [Fig. 2(b)] enables a fundamentally different electrically controlled transport behavior compared to its weak counterpart. In this case, we note that the resonant conditions in Eqs. (23)–(III.2) still apply. In particular, the spin polarization of the Fermi surface in a strong AM can be switched by tuning the energy, as illustrated in Fig. 3(a-iv), which is a behavior absent in weak AMs. As a result, spin-down (spin-up) electrons with $\theta_{k}^{-}<\theta_{c}^{-}$ ($\theta_{k}^{+}<\theta_{c}^{+}$) are blocked for $E>U_{g}$ ($E<U_{g}$), as shown in Fig. 3(c-ii) and Fig. 3(c-iii), respectively. This can also be understood from the possible Fermi surface contours at positive energy $E$, illustrated in the inset of Fig. 2(b). In this regime, spin-up (spin-down) states are absent when $a|k_{x}|<\sqrt{(E-U_{g})/(t+J)}$ ($a|k_{y}|<\sqrt{(E-U_{g})/(t+J)}$). Conversely, at negative energies, spin-up (spin-down) states vanish when $a|k_{y}|<\sqrt{(E-U_{g})/(t-J)}$ ($a|k_{x}|<\sqrt{(E-U_{g})/(t-J)}$), see Fig. 3(a-iv).
Consequently, transmission at $E>U_{g}$ ($E<U_{g}$) is dominated by spin-up (spin-down) electrons, as shown in Fig. 3(c-i)-(c-iv). This feature is most pronounced for normally incident electrons, which is exclusively carried by spin-up (spin-down) states for $E>U_{g}$ ($E<U_{g}$). Spin-up electrons with $\theta_{k}\neq 0$ are blocked for $E<U_{g}$ due to the large forbidden angle range, leading to a fully spin-polarized transmission and thus conductance [Eq. (20)] in this energy regime, while for $E>U_{g}$, spin-down electrons with $\theta_{k}>\theta_{c}^{-}$ can still contribute to transmission at $E>U_{g}$, thereby reducing the degree of spin polarization. This behavior is reflected in the conductance and spin polarization shown in Fig. 3(c-v) and Fig. 3(c-vi).
III.4 Effects of the altermagnetic orientation $\theta_{J}$
We now examine the influence of the altermagnetic orientation angle $\theta_{J}$ on the resulting spin polarization. For a $d_{x^{2}-y^{2}}$-wave weak AM with $\theta_{J}=0$, our analysis reveals that spin-polarized transport favors the spin-down channel, yielding $P<0$, tunable via the incident electron energy $E$ and the barrier height $V$ [see Fig. 3(b-vi)]. Rotating the orientation to $\theta_{J}=\pi/2$, corresponding to an equivalent $d_{x^{2}-y^{2}}$-wave AM rotated by $90^{\circ}$, reverses the spin polarization, resulting in $P>0$. This configuration can be interpreted as reorienting the AM junction along the $y$-direction rather than the $x$-direction in Fig. 3. The reversal spin polarization arises from the interchange of spin-split subbands, effectively mapping the spin index $\sigma\rightarrow-\sigma$ in key transport signatures such as resonant energy levels and critical incident angles, thereby inverting the sign of $P$. In contrast, for a $d_{xy}$-wave AM with $\theta_{J}=\pi/4$, transport is dominated by spin-degenerate propagation along the nodal directions ($y=\pm x$), resulting in spin-independent transmission and vanishing net spin polarization.
Orientation-dependent transport reveals their alternating behavior of AM between ferromagnetic and antiferromagnetic characteristics. Transport along directions of maximal spin splitting (e.g., $\theta_{J}=0$ or $\pi/2$) exhibits ferromagnetic-like features, whereas spin-degenerate directions (e.g., $\theta_{J}=\pi/4$) give rise to antiferromagnetic-like effects. This anisotropy leads to orientation-dependent suppression of Andreev reflection [50, 51, 53, 52], the emergence of $\pi$-state Josephson currents, and nonreciprocal diode effects [57, 58, 59, 60], all without the need for an external magnetic field. The intrinsic spin splitting in $d_{x^{2}-y^{2}}$-wave AM thus provides a highly tunable spin-selective transport, crucial for realizing spintronic functionalities in systems with zero net magnetization.
IV Gate-Controlled Spin Filtering and Spin Valve Effect in Altermagnet junctions
The distinct energy dependence of the spin-resolved transmission and the resulting conductance (Fig. 3), enables gate-tunable spin transport in AM-based junctions, offering a promising platform for spintronic applications without net magnetization. In this section, we propose two electrically controlled spintronic devices that utilize the anisotropic spin splitting in $d_{x^{2}-y^{2}}$-wave AMs: a spin filter [Fig. 3(a) and Fig. 4] and a spin valve [Fig. 5]. In both cases, the spin-resolved conductance $G_{\sigma}$ and spin polarization $P$ can be modulated by gate voltages applied within the AM region and at the tunneling barrier at the AM–lead interface. Notably, by exploiting the opposite band curvature of the spin-split subbands in strong AMs, we demonstrate a double-gate-controlled spin valve. This device exhibits a lower (higher) conductance state when the two gates are opposite (aligned), analogous to conventional spin valves based on ferromagnetic bilayers [12, 13], but realized here in a system with zero net magnetization.
IV.1 Gate-controlled spin filter effects
The distinct energy-dependent conductance and polarization of strong and weak AMs enable gate- and barrier-controlled generation of spin-polarized conductance. To show this effect, Figs. 4 display the spin-resolved conductance and spin polarization of the total conductance as functions of the gate voltage $U_{g}$.
In the weak AM regime, the conductance exhibits spin-dependent oscillations as the gate potential $U_{g}$ increases from zero [Fig. 4(a)], arising from quantum confinement effects [77, 78]. The conductance peaks appear when the resonance condition [Eq. (23)] is satisfied. Between the resonance peaks, one spin channel is strongly suppressed by the tunneling barrier, while the other remains finite, leading to asymmetric spin-resolved conductances. This asymmetry gives rise to finite spin polarization $P$ [Fig. 4(b)]. At low barrier strength (e.g., $V=0$), the conductance is only partially spin-polarized, whereas at higher barriers (e.g., $V=10$), a fully spin-polarized conductance is realized with a switchable polarization between $P=+1$ and $P=-1$ by tuning the gate voltage $U_{g}$.
In the case of strong AM, the conductance as a function of $U_{g}$ exhibits a highly tunable behavior, persisting even for $U_{g}<0$, see Fig. 4(c). This can be understood by the unique Fermi surface geometry of strong AM [Fig. 3(a-iv)], which leads to a regime in which the conductance is exclusively carried by spin-up electrons for $U_{g}>0$ and by spin-down electrons for $U_{g}<0$. As a result, the spin polarization reaches $P=\pm 1$ and can be flipped simply by tuning the gate voltage, as shown in Fig. 4(d). Importantly, this spin filtering effect remains robust regardless of the interface transparency set by the tunneling barrier $V$, indicating that it arises from intrinsic bulk properties of the strong AM, rather than interfacial effects in the junction.
Figure 4: Gate-controlled spin-filter effect for (upper panel) weak and (lower panel) strong altermagnets, corresponding to Fig. 3. (a) and (c): Spin-resolved conductance as a function of gate voltage $U_{g}$ at fixed tunneling barrier $V=5$. (b) and (d): Spin polarization as a function of $U_{g}$ for various values of $V$. The Fermi level is set to $E=0.1t$, and all other parameters are identical to those in Fig. 3.
Beyond the demonstrated control of spin polarization, the AM-based spin filter offers several key advantages for practical implementation. First, because the spin polarization originates from intrinsic Fermi surface geometry rather than interface interference or external magnetic fields, the spin filtering mechanism is expected to be robust against moderate disorder and thermal broadening [12, 13, 14, 15, 16]. Second, the ability to reversibly and continuously tune both the magnitude and sign of $P$ using a single electrostatic gate, without altering material composition or invoking spin–orbit coupling, distinguishes this approach from conventional spintronic platforms [85, 86]. Finally, the momentum-selective spin filtering effect in strong AMs may be leveraged to design functional interfaces with superconductors or topological materials [61, 49, 62, 63, 64, 65] and spin-sensitive Josephson effects [46, 52, 59, 60, 56, 57, 58]. These features collectively underline the strong AM as a versatile and scalable building block for spintronic devices.
IV.2 Electrically controlled spin valve based on strong AMs
The gate-controlled spin polarization flipping in the strong-AM spin filter can be further extended to realize a fully electrically tunable spin valve, a functionality inaccessible in weak AMs considered in this work. A conventional spin valve consists of two spin filters connected in series [13, 12]. Each spin filtering effect is composed of a magnetic layer, where only electrons with spin aligned to the magnetization contribute to the current. High (low) conductance state occurs when the magnetization directions of the two ferromagnetic layers are aligned (opposite), denoted as parallel (antiparallel) configurations, which is switched typically by external magnetic field [12, 13, 14, 15, 16].
The gate-tunability in the strong-AM junction enables the realization and manipulation of a spin valve effect with switchable parallel-like and antiparallel-like configurations without net magnetization and external magnetic field. The proposed device consists of a strong AM junction with two independently gated regions [Fig. 5(a)], characterized by gate voltages $U_{g1}$ and $U_{g2}$. Each gated region serves as a spin filter with spin-up (spin-down) polarization realized by the positive (negative) gate [Figs. 4(c) and (d)]. As a result, when both gates are applied in the same direction ($U_{g1}U_{g2}>0$), spin-polarized conducting channels of identical spin species are activated in both regions [Fig. 5(a)], analogous to a conventional spin valve in the parallel configuration. Conversely, when the gate voltages are of opposite sign ($U_{g1}U_{g2}<0$), different spin-polarized channels are activated in each region [Fig. 5(a)], mimicking the antiparallel configuration of a conventional magnetic spin valve.
Crucially, unlike spin valves composed of ferromagnetic bilayers, where parallel-antiparallel switching requires external magnetic fields to reorient the magnetizations [12], the strong-AM spin valve we proposed is entirely dependent on electrostatic gating. The effective spin polarization is governed by the intrinsic momentum-dependent spin-splitting and open Fermi surface of the strong AM [Fig. 2(b)], without requiring net magnetization or magnetic field control. This gate-controlled analog of ferromagnetic bilayers establishes strong AMs as a compelling platform for a fully electrical spintronic device.
Figure 5: (a) Schematic illustration of a double-gate-controlled spin valve based on a strong altermagnetic heterostructure. The configuration enables gate-tunable switching between parallel ($U_{\text{g}1}U_{\text{g}2}>0$) and antiparallel ($U_{\text{g}1}U_{\text{g}2}<0$) regimes. (b) Total conductance and (c) spin polarization as functions of the gate voltages $U_{\text{g}1}$ and $U_{\text{g}2}$. The length of the altermagnetic region is $d=50a$, and the tunneling barrier is set to $V=0$.
To explain the transport mechanism within this electrically modulated spin valve, we use Eq. (III.1) but with an electrostatic potential in AM given by
$U_{g}\Theta\left(+x\right)\Theta\left(d-x\right)$
$\rightarrow$ $U_{g1}\Theta\left(+x\right)\Theta\left(d/2-x\right)+U_{g2}\Theta\left(x-d/2\right)\Theta\left(d-x\right)\text{,}$
where, for simplicity, we assume that each gated region occupies half of the junction length $d$. The spin-resolved transmission $T_{\sigma}$, total conductance $G$, and spin polarization $P$ can be calculated by solving the quantum scattering problem using the scattering states
$\psi^{\sigma}(x)=\left\{\begin{array}[]{ll}e^{ik_{0}x}+r_{\sigma}e^{-ik_{0}x},&x\leq 0,\\[4.0pt] c_{1}e^{ik_{1+}^{\sigma}x}+c_{2}e^{ik_{1-}^{\sigma}x},&0<x\leq d/2,\\[4.0pt] c_{3}e^{ik_{2+}^{\sigma}x}+c_{4}e^{ik_{2-}^{\sigma}x},&d/2<x\leq d,\\[4.0pt] t_{\sigma}e^{ik_{0}x},&x>d,\end{array}\right.$ (28)
with
$k_{1,2\pm}^{\sigma}=k_{\pm}^{\sigma}(U_{g}\rightarrow U_{g1,2}).$ (29)
In addition to the boundary conditions in Eq. (16), the wavefunction must also satisfy
$\Psi^{\sigma}\left[(d/2)^{+}\right]$ $=$ $\Psi^{\sigma}\left[(d/2)^{-}\right],$ (30)
$\partial_{x}\Psi^{\sigma}\left[(d/2)^{+}\right]$ $=$ $\partial_{x}\Psi^{\sigma}\left[(d/2)^{-}\right].$
By solving the scattering problem using the same procedure as in Sec. III.1, the spin-resolved transmission $T_{\sigma}$ in the spin valve configuration is obtained. Using Eqs. (20)–(22), the total conductance $G$ and the resulting spin polarization $P$ are evaluated. Their dependence on the gate voltages $U_{g1}$ and $U_{g2}$ is shown in Figs. 5(b) and 5(c), respectively.
Figure 5(b) demonstrates the switching between high and low conductance states by tuning the relative values of the gate voltages $U_{g1}$ and $U_{g2}$ in the spin valve based on a strong AMs [Fig. 5(a)]. This effect originates from the opposite spin-dependent band-opening directions in strong AMs [Fig. 5(a)]. As a result, transmission through each gated region is dominated by spin-up (spin-down) electrons for positive (negative) gate voltages, consistent with the spin-filtering behavior shown in Fig. 4(d). A high-conductance state arises when $U_{g1}U_{g2}>0$, indicating aligned spin polarizations in the two gated AM regions [Fig. 5(b)]. This configuration is analogous to the parallel alignment of magnetizations in conventional ferromagnetic spin valves [13, 14]. More specifically, for $U_{g1},U_{g2}>0$ ($<0$), transport is fully carried by spin-up (spin-down) carriers, yielding a spin polarization of $P=+1$ ($-1$), as shown in Fig. 5(c). In contrast, when $U_{g1}U_{g2}<0$, the opposite spin polarization in the two gated regions leads to a strong suppression of conductance due to spin-filter mismatch, resulting in an ill-defined $P$ through Eq. (22). This behavior is reminiscent of the antiparallel configuration in conventional spin valves, where spin-polarized electrons are blocked by the misaligned magnetic orientation of the second layer [13, 14].
The gate-tunable switching between high and low conductance states in the strong-AM-based spin valve originates from the opposite spin-dependent band-opening directions [Fig. 5(a)], which is absent in its weak-AM counterpart. This functionality arises from electrically controlled spin polarization in each gated strong-AM region, analogous to the spin-filtering behavior in conventional spin valves composed of ferromagnetic bilayers. However, unlike conventional spin valves, where switching between parallel and antiparallel configurations typically requires external magnetic fields or spin-transfer torques [13, 14], our design operates entirely through electrostatic gating. This enables faster, more energy-efficient control and facilitates integration with standard semiconductor platforms [69]. Moreover, the pronounced conductance contrast between the two gate configurations, together with the ability to reversibly switch spin polarization, offers a promising route toward nonvolatile spin-based logic and memory elements [12, 13, 14, 15, 16]. These results underscore the potential of strong AM as a robust platform for scalable, field-free, and all-electric spintronic device architectures without net magnetization.
As a final remark, although the double-gate spin valve is inaccessible in the weak-AM-based junction in the system concerned here. In recently developed ferroelectric-switchable AMs [71, 72], the spin-split Fermi surface can be reoriented by reversing the ferroelectric polarization. This effect is analogous to an electric-field-induced rotation of the altermagnetic orientation angle from $\theta_{J}=0$ to $\theta_{J}=\pi/2$ in the present model. A gate-controlled spin valve is also conceivable in weak-AM-based junctions. Although such systems often involve multiple bands, which lie beyond the scope of the two-band model considered in this work, they also demonstrate the potential spintronics application of the AM-based junctions.
V Conclusions
We have investigated spin-selective transport in heterostructures composed of altermagnetic materials and normal-metal leads, focusing on the contrasting behaviors between weak and strong altermagnets. By analyzing the interplay between altermagnetic exchange interaction and kinetic energy, we classified altermagnets into weak and strong regimes, each exhibiting distinct spin-resolved Fermi surface configurations. These differences manifest in qualitatively distinct transport signatures. While weak altermagnets support partial spin filtering via resonant tunneling enhanced by interface barriers, strong altermagnets intrinsically support fully spin-polarized conductance due to the absence of one spin species near the Fermi surface. Building upon these findings, we proposed an electrically tunable spin valve based on strong altermagnets. In this device, double-gate control allows transitions between parallel and antiparallel spin configurations without requiring magnetic fields or net magnetization, unlike conventional ferromagnetic spin valves. This functionality stems from the nonrelativistic spin-splitting inherent to strong AMs and highlights their potential as a platform for scalable, all-electrical spintronic devices. Our results demonstrate that altermagnets, particularly in the strong regime, offer a robust and versatile route to realizing spintronic functionalities through purely electric control. This work paves the way for future experimental efforts in designing gate-tunable spin filters and spin valves in candidate materials such as weak altermagnets RuO2 [40, 23, 25] and strong altermagnets Mn5Si3 [26, 42].
Acknowledgements.
P.-H. Fu and Q. Lv contribute equally to this work. P.-H. Fu appreciates the support from W. Xu and the discussion from K. W. Lee and Z. Yuan. Y. Xu acknowledges financial support from the Scientific Research Starting Foundation of Ningbo University of Technology (Grant No. 2022KQ51). J. C. acknowledges financial support from the Carl Trygger’s Foundation (Grant No. 22: 2093), the Sweden-Japan Foundation (Grant No. BA24-0003), the Göran Gustafsson Foundation (Grant No. 2216), and the Swedish Research Council (Vetenskapsrådet Grant No. 2021-04121). J.-F. L. acknowledges financial support from the National Natural Science Foundation of China (Grant No. 12174077) and the Joint Fund with Guangzhou Municipality under Grant No. 202201020238. X.-L.Yu acknowledges financial support from the Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023A1515011852).